# Truthful Text Sanitization Guided by Inference Attacks

### v1

@@ Abstract
The purpose of text sanitization is to rewrite those text spans in a document that may directly or indirectly identify an individual, to ensure they no longer disclose personal information. Text sanitization must strike a balance between preventing the leakage of personal information (privacy protection) while also retaining as much of the document’s original content as possible (utility preservation). We present an automated text sanitization strategy based on generalizations, which are more abstract (but still informative) terms that subsume the semantic content of the original text spans. The approach relies on instruction-tuned large language models (LLMs) and is divided into two stages. The LLM is first applied to obtain truth-preserving replacement candidates and rank them according to their abstraction level. Those candidates are then evaluated for their ability to protect privacy by conducting inference attacks with the LLM. Finally, the system selects the most informative replacement shown to be resistant to those attacks. As a consequence of this two-stage process, the chosen replacements effectively balance utility and privacy. We also present novel metrics to automatically evaluate these two aspects without the need to manually annotate data. Empirical results on the Text Anonymization Benchmark show that the proposed approach leads to enhanced utility, with only a marginal increase in the risk of re-identifying protected individuals compared to fully suppressing the original information. Furthermore, the selected replacements are shown to be more truth-preserving and abstractive than previous methods.
@@
journal: Information Sciences
@@
[inst2]organization=Department of Computer Engineering and Mathematics. CYBERCAT.,addressline=Universitat Rovira i Virgili, city=Tarragona, postcode=43007, country=Spain
@@ 1 Introduction
Text documents play a key role in many sectors, such as healthcare, legal services or public administrations. However, those documents often contain personal data that, if released without adequate protection, can pose significant privacy risks to the individuals they concern. Regulations such as the General Data Protection Regulation (GDPR) [15] impose strict constraints how personal information should be protected. In particular, such data cannot be shared with third parties or employed for secondary purposes without a proper legal ground, such as the explicit consent of the individuals the data refers to. However, if the data can be fully anonymized, those restrictions no longer apply, as the resulting data is by definition no longer considered personal. Many approaches have been proposed to perform this anonymization [24], although their compliance relative to legal frameworks such as GDPR remains unclear [44].
Text sanitization is a type of text anonymization technique that involves i) detecting and ii) protecting sensitive text spans. Most approaches in the existing literature focus on the detection step, and employ straightforward methods to protect the detected text spans, such as suppressing them or replacing them by generic semantic categories indicating named entity types such as PERSON or LOCATION [21, 46, 47, 18, 25, 9, 30]. However, this relatively crude level of protection can undermine the usability of the protected data for research and analysis, which is the ultimate goal of data sharing. As highlighted in the data privacy literature [19], data protection should aim at balancing the trade-off between the attained level of privacy and the preservation of data utility. For textual data, this means designing effective text sanitization mechanisms that preserve the meaning of the original text as much as possible while fulfilling the privacy needs of the data release.
@@ 1.1 Contributions and plan
Large Language Models (LLMs) are particularly effective for text rewriting tasks, among others, due to their ability to understand both the detailed and broader contextual aspects of language. This enables them to rephrase content while preserving its intended meaning, making them valuable for applications like text simplification [22] and editing [41]. Although LLMs can be adapted for text rewriting through fine-tuning on task-specific data, a simpler approach is to rely on instruction-tuned LLMs to directly prompt the model with rewriting instructions.
@@
Unlike previous privacy-oriented text rewriting methods [5, 13], our method adopts a two-stage strategy. Given a text span to replace, we first (1) generate a sorted list of generalization candidates at different abstraction levels, and then (2) select the most suitable replacement among those alternatives. Crucially, this selection of the best replacement is done by testing whether the LLM can guess the original span from the replacement and the rest of the text (inference attack). As LLMs have been shown to be effective adversaries in anonymization tasks in previous studies [31], leveraging their predictions is a crucial step towards ensuring the privacy of the generated replacements, particularly truthful ones. Separating these steps not only allows balancing and customizing the privacy-utility trade-off, but it also enables a more granular evaluation of the rewriting process.
We evaluate and compare our method against several baselines on the Text Anonymization Benchmark (TAB) [32], a dataset explicitly tailored for the evaluation of text sanitization methods. As text sanitization is by definition a process that involves the processing of (sometimes highly sensitive) personal data, our approach relies on the use of open-weights models, as those can be run locally without necessitating the transfer of data to third-party cloud solutions. Empirical results show that our method achieves the best privacy-utility trade-off, while keeping the protected data truthful to their original content. We also show how privacy and utility can be measured without reference annotations.
@@ 2 Background and related works
Text anonymization aims to remove or mask any personally identifiable information (PII) that may lead to re-identifying the subject(s) referenced in the text and/or disclose their confidential information.
@@
For detecting PII, the vast majority of automatic text anonymization approaches rely on Named Entity Recognition (NER) [21, 46, 47, 18, 25, 9], which assume that named entities (NEs), due to their specificity, may facilitate re-identification and should therefore be subjected to protection. However, because not all PIIs fit within the limited set of NE types supported by NER tools, NER-based methods may miss a significant amount of disclosive terms [24]. To mitigate this issue, other approaches grounded on the broader notion of disclosure risk have been proposed [17, 30, 35].
The above approaches provide solutions for the detection of disclosive terms, but do not specify how those terms ought to be edited in the document. Two editing strategies are possible. Whereas document redaction directly suppresses or blacks out terms, thereby maximizing privacy at the cost of the utility and readability of the protected documents, text sanitization replaces terms with broader, more abstract generalizations [8]. The latter achieves a better privacy-utility trade-off while preserving the truthfulness of the information, as generalizations, by definition, should retain the core semantic content of the original terms.
@@
The majority of current text anonymization methods use either strict redaction or replace the disclosive terms by their corresponding NE category (such as PERSON or LOCATION), which is a basic form of sanitization [24]. Both strategies retain little to no meaningful information from the original terms. More sophisticated approaches leverage structured knowledge bases, such as ontologies, or, more recently, language models, to retrieve potential replacements. We survey those methods in the following subsections.
2.1 Sanitization from Knowledge Bases
Automatic text sanitization methods have traditionally leveraged structured knowledge bases [10, 3, 2, 37, 38, 36, 17, 29], such as WordNet, the now defunct Open Directory Project (ODP) or YAGO, to retrieve term generalizations to be used as replacements.
@@ 2.1 Sanitization from Knowledge Bases
They propose different mechanisms to select the most appropriate generalization from a list of candidates. In [17], the authors leverage the Word2Vec [28] word embedding model to identify terms that are semantically related to the individual to be protected. If a term’s embedding has a cosine similarity with the individual’s name above a user-defined threshold, it is considered disclosive, and it is replaced by its most specific generalization (retrieved from YAGO) whose cosine similarity falls below the threshold. In [37, 36] the authors employ a similar approach, but this time selecting the most appropriate replacement through measures of the term informativeness and mutual information. [3] propose to select generalizations so that they encompass a number of specializations (or hyponyms) that would result in $t$ or more plausible documents. [2] employs a contingency table that quantifies the degree of correlation between each pair of textual terms and a taxonomy modeling term generalizations. Finally, [29] rely on human judgments to select the most appropriate generalization for each sensitive term based on its context.
@@
Nonetheless, sanitization methods based on “static” knowledge bases have limited applicability due to their relatively narrow coverage of disclosive terms. This limited coverage is particularly salient for named entities, which are rarely present in such knowledge bases, while they are often one of the most crucial type of terms to sanitize. Moreover, the largest knowledge bases, such as YAGO or ODP may not provide consistent generalizations, as they originate from arbitrary categories, labels and descriptions (such as Wikipedia categories) that do not fulfill the subsumption property; and this may hamper the truthfulness of the sanitized outcomes.
2.2 Sanitization with Language Models
@@
More recently, a number of studies have also explored LLMs for text anonymization, both for detecting and for protecting sensitive text spans. Such models successfully overcome the aforementioned coverage issues of knowledge bases as they can (1) generate replacements for any input span and (2) can ensure those replacements can be inserted in the document in a grammatically meaningful manner.
The MAPA anonymization toolkit [5] is an open source tool capable of both detecting a variety of PII types and replacing them with alternatives sampled from the most likely next tokens based on BERT language models fine-tuned for a variety of European languages and domains including administration, health and law. [1] relies on the same type of models, but without fine-tuning for PII replacement. Those two approaches, however, generate replacements that are restricted to the same token length as the original sensitive term. This restriction has two important shortcomings. The first is that it rules out replacements that are either shorter or longer (in terms of number of tokens) than the original text span. Furthermore, this restriction may be easily exploited by an adversary seeking to infer the original content of the replaced span.
@@
More recently, larger, instruction fine-tuned language models such as those offered by ChatGPT [45] have also been proposed for protecting sensitive information. Instead of fine-tuning, a common way to get these models to perform a task is via an instruction (prompt) and optional examples. Such prompting methods are also integrated in existing de-identification tools, such as Microsoft Presidio. Given a text, the PII detected with the Presidio pipeline can be replaced with GPT 3.5-generated alternatives in a single step [^2] https://github.com/microsoft/presidio/blob/main/docs/samples/python/synth_data_with_openai.ipynb.
@@
None of the LLM-based works mentioned above focus explicitly on generating truthful replacements – that is, replacements that subsume the semantic content of the original text span. As a result, the selected replacement may significantly distort the original terms’ meaning. A study in this direction on a closely related tasks is described in [13], who look at abstracting text spans containing self-disclosure, “the communication of personal information to others”. The categories of personal information employed in this study are a fine-grained version of those typically considered by previous anonymization approaches, and mostly fall under demographic attributes. The sensitive text spans in [13], however, are often phrases and clauses (written in first person) rather than words or named entities. Their proposed task can be thus viewed as a specialized case of protecting sensitive information. PII protection is performed step-wise, where self-disclosure spans are first detected with a BERT model fine-tuned on a manually annotated dataset. The text spans expressing personal information are then re-written with an open-weights instruct LLM (Llama-2-7B) fine-tuned on a dataset with replacements automatically generated using GPT-3.5 and GPT-4.
@@
In contrast, our work leverages open-weights instruct LLMs without additional fine-tuning to iteratively select truthful generalizations from a set of alternatives with different abstraction levels, using inference attacks to test whether a given replacement effectively conceals the initial content of the span.
@@ 2.3 Evaluation
A standard evaluation strategy for text anonymization methods is to contrast their outcomes against manual annotations, which are considered as ground truths. The anonymization quality is quantified via the precision and recall metrics, which are used as proxies of the preserved utility and the attained privacy protection, respectively [24, 32, 35, 16, 21].
@@
However, these evaluation metrics are only applicable to assess the quality of the detection step, and are unable to assess the suitability of the replacements used for sanitization. Indeed, evaluating sanitized texts against a manually constructed ground truth is challenging because i) there can be multiple valid (i.e., non-disclosive) combinations of replacements for a given text and, ii) replacement candidates are, quite often, unbounded. Establishing a single ground truth for the evaluation is thus impossible. Moreover, requesting human annotators to provide privacy-preserving replacements (or, at the very least, chose appropriate replacements for predefined lists of candidates) requires even more effort than the –already quite costly– detection of disclosive text spans [29]. As a result, the few sanitization strategies proposed in the literature are largely untested [10, 20, 2, 37, 38, 36, 17, 29].

### v2

@@ Abstract
Text sanitization aims to rewrite parts of a document to prevent disclosure of personal information. The central challenge of text sanitization is to strike a balance between privacy protection (avoiding the leakage of personal information) and utility preservation (retaining as much as possible of the document’s original content). To this end, we introduce a novel text sanitization method based on generalizations, that is, broader but still informative terms that subsume the semantic content of the original text spans. The approach relies on the use of instruction-tuned large language models (LLMs) and is divided into two stages. Given a document including text spans expressing personally identifiable information (PII), the LLM is first applied to obtain truth-preserving replacement candidates for each text span and rank those according to their abstraction level. Those candidates are then evaluated for their ability to protect privacy by conducting inference attacks with the LLM. Finally, the system selects the most informative replacement candidate shown to be resistant to those attacks. This two-stage process produces replacements that effectively balance privacy and utility.
We also present novel metrics to evaluate these two aspects without needing to manually annotate documents. Results on the Text Anonymization Benchmark show that the proposed approach, implemented with Mistral 7B Instruct, leads to enhanced utility, with only a marginal (<1<1 p.p.) increase in re-identification risk compared to fully suppressing the original spans. Furthermore, our approach is shown to be more truth-preserving than existing methods such as Microsoft Presidio’s synthetic replacements.
@@
journal: Applied Soft Computing
@@
[inst2]organization=Department of Computer Engineering and Mathematics. CYBERCAT,addressline=Universitat Rovira i Virgili, city=Tarragona, postcode=43007, country=Spain
@@ 1 Introduction
Text documents play a key role in many sectors, including healthcare, legal services and public administrations. However, those documents often contain personal data that, if released without adequate protection, can pose significant privacy risks to the individuals they concern. Regulations such as the General Data Protection Regulation (GDPR) [24] impose strict constraints on how personal information should be protected. In particular, personal information cannot be shared with third parties or employed for secondary purposes without a proper legal ground, such as the explicit consent of the individuals to whom the data refers.
One strategy to overcome those constraints is to anonymize the data, in which case the data is no longer considered personal, and thus outside the scope of those regulations [24]. Beyond regulatory compliance, anonymization also helps balance the pursuit of data-driven research and innovation with the protection of individual rights. By reducing the risk of re-identification, anonymization fosters greater trust between individuals and institutions, encouraging more open data sharing practices that ultimately benefit society as a whole.
Multiple text anonymization techniques have been proposed in the literature [36]. In particular, text sanitization prevents the disclosure of personal information by: (i) detecting spans that contain personally identifiable information (PII), and (ii) masking those spans. Most works have focused on the first step (PII detection), and very often rely on simple masking techniques, such as removing PIIs or replacing them by generic semantic categories such as PERSON or LOCATION) [32, 64, 65, 29, 37, 10, 44]. However, this relatively crude level of protection can undermine the downstream utility of the sanitized texts, which is the ultimate goal of data sharing. As the data privacy literature has highlighted [30], data protection should aim at balancing the trade-off between the attained level of privacy and the preservation of data utility. For textual data, this means designing effective text sanitization mechanisms that preserve the original text’s meaning as much as possible while satisfying the privacy needs of the data release.
In this context, large language models (LLMs), which have shown strong capabilities in a wide range of natural language understanding and generation tasks, have emerged as promising tools for enhancing text sanitization. Recent studies have demonstrated their effectiveness in text editing [58] and simplification [33] with strong zero- and few-shot performance due to instruction tuning [42].
Despite these advances, current text sanitization approaches rarely leverage LLMs to generate replacements for detected PII spans. Most existing methods rely on suppression or named entity-based substitutions, which fail to produce coherent and realistic text. As a result, there is currently no principled method for generating truthful and utility-preserving replacements of sensitive text spans. Addressing this gap is essential for enabling safe and effective data sharing under privacy regulations while maintaining the semantic integrity and analytical value of textual data.
@@ 1.1 Contributions and plan
Our work brings the following contributions:
  • 1.
    Unlike previous privacy-oriented text rewriting methods [6, 18], the presented method adopts a two-stage strategy. Given a text span to replace, we first (1) generate a sorted list of generalization candidates at different abstraction levels, and then (2) select the most suitable replacement among those alternatives. This selection of the best replacement is done by testing whether the LLM can guess the original span from the replacement and the rest of the text (inference attacks). LLMs have indeed shown to be effective adversaries in anonymization tasks in previous studies, either on their own [45] or augmented with a retrieval component [12]. The combination of those two steps not only allows for balancing and customizing the privacy-utility trade-off, but also enables a more granular evaluation of the rewriting process.
  • 2.
    We present an evaluation framework for text sanitization that measures privacy and utility without reference annotations. This framework includes a novel utility metric, Text Preserved Similarity (TPS), which measures the similarity between original and sanitized documents according to both information content and semantic similarity.
  • 3.
    We evaluate and compare our method against several baselines on the Text Anonymization Benchmark (TAB) [46], a dataset explicitly tailored for the evaluation of text sanitization methods. Empirical results show that our method achieves the best privacy-utility trade-off, while keeping the protected data truthful to their original content.
As text sanitization is by definition a process that involves the processing of (sometimes highly sensitive) personal data, our approach relies on the use of open-weights LLMs, as those can be run locally without necessitating the transfer of data to third-party cloud solutions.
@@ 2 Background and related works
2.1 Problem overview
Text anonymization aims to remove or mask personally identifiable information (PII) that may lead to re-identifying the subject(s) referenced in the text and/or disclose their confidential information. This process should consider all PII, including both directly identifying information –that is, information that refers to single subjects, such as full names or passport numbers–, and quasi-identifying terms which do not enable re-identification when considered in isolation, but may do so when combined with other quasi-identifiers, such as the combination of gender, birth date and postal code.
@@ 2.1 Problem overview
For detecting PII, the majority of automatic text anonymization approaches rely on Named Entity Recognition (NER) [32, 64, 65, 29, 37, 10]. Indeed, due to their specificity, named entities (NEs) often increase the re-identification risk. However, because not all PII fit within the limited set of NE types supported by NER tools, NER-based methods may miss a significant amount of disclosive terms [36]. To mitigate this issue, other detection approaches grounded on the broader notion of disclosure risk have been proposed [27, 44, 51]. In parallel, some methods based on differential privacy [19] have been proposed for text protection [22]. These typically treat texts as bags-of-words, replacing sensitive terms via the exponential mechanism or adding Laplace noise to term counts. However, they often degrade utility and readability, failing to preserve truthfulness and document structure.
Once disclosive terms are detected in a first processing stage, two editing strategies are possible. Whereas document redaction directly suppresses or blacks out terms, thereby maximizing privacy at the cost of the utility and readability of the protected documents, text sanitization replaces PII spans with broader, more abstract terms [9]. By obscuring the disclosive information while preserving as much as possible of the original text’s semantic content and readability, text sanitization seeks to offer a better privacy-utility trade-off than redaction. The choice of replacement strategy plays a crucial role in this balance. Generalizations, which abstract sensitive terms without distorting their meaning, help maintain the truthfulness of the document and are particularly valuable in domains like medicine or law, where accuracy is essential. In contrast, substituting terms with unrelated or misleading ones (e.g., replacing broken arm with bird flu) can compromise the integrity of the information and lead to serious misinterpretations.
2.2 Sanitization
@@
The majority of text anonymization methods use either strict redaction or replace the disclosive terms by the label of their corresponding category, which is a basic form of sanitization [36]. Both strategies retain little to no meaningful information from the original terms. More sophisticated approaches leverage structured knowledge bases, such as ontologies, or, more recently, language models, to retrieve potential replacements. We survey those methods below.
2.2.1 NER-based sanitization
Following the broad usage of NER for PII detection, the most common practice for replacements is to use the named entity labels (e.g., PERSON or LOCATION). This straightforward approach maintains truthfulness to the extent that the classification is accurate, because the labels themselves offer a coarse generalization of the original terms. A notable exception is the MISC (miscellaneous) category [61], which is used in some NER systems to cover entities that do not fit into more specific categories. Replacing such entities with MISC is effectively equivalent to redaction, as it conveys no semantic information about the original term. Overall, NER-based sanitization offers limited utility preservation, as the original content of the PII spans is mostly lost and the readability of the text is severely reduced.
2.2.2 Sanitization from knowledge bases
@@
Other text sanitization methods have leveraged structured knowledge bases [14, 4, 3, 53, 54, 52, 27, 43], such as WordNet, the now defunct Open Directory Project (ODP) or YAGO, to retrieve term hypernyms (i.e., generalizations) to be used as replacements. Those methods rely on dedicated mechanisms to select the most appropriate generalization from a list of candidates. In [27], the authors leverage the Word2Vec [41] word embedding model to identify terms that are semantically related to the individual to be protected. If a term’s embedding has a cosine similarity with the individual’s name above a user-defined threshold, it is considered disclosive, and it is replaced by its most specific generalization (retrieved from YAGO) whose cosine similarity falls below the threshold. In [53, 52] the authors employ a similar approach, but this time selecting the most appropriate replacement through measures of the term informativeness and mutual information. [4] propose to select generalizations so that they encompass a number of specializations (or hyponyms) that would result in $t$ or more plausible documents. [3] employs a contingency table that quantifies the degree of correlation between each pair of textual terms and a taxonomy modeling term generalizations. Finally, [43] rely on human judgments to select the most appropriate generalization for each sensitive term based on its context.
@@
Nonetheless, sanitization methods based on “static” knowledge bases have limited applicability due to their relatively narrow coverage of disclosive terms. This limited coverage is particularly salient for named entities, which are rarely present in such knowledge bases, while they are often one of the most crucial type of terms to sanitize. Moreover, the largest knowledge bases, such as YAGO or ODP may not provide consistent generalizations, as they originate from arbitrary categories, labels and descriptions (such as Wikipedia categories) that do not systematically satisfy the subsumption property – that is, the requirement that a generalized term should denote a concept that logically encompasses all instances of the original text span.
2.2.3 Sanitization with language models
@@
A number of recent studies have also explored the use of language models (LMs) for text anonymization, both for detecting and for protecting sensitive text spans. Such models successfully overcome the aforementioned coverage issues of knowledge bases, as they can (1) generate replacements for any input span and (2) ensure those replacements integrate seamlessly into the surrounding text, preserving grammaticality and fluency.
The MAPA anonymization toolkit [6] is an open source tool that supports the detection and replacement of a variety of PII types. The replacements are sampled from the most likely tokens based on BERT language models fine-tuned for a variety of European languages and domains including administration, health and law. [2] rely on similar models, but without fine-tuning for PII replacement. In both cases, the tokens of the sensitive terms are initially replaced by the special token [MASK]. The BERT model then predicts for each mask a distribution over possible tokens based on the context. The replacements are finally sampled from these distributions. Due to the use of small-size LMs, those approaches only require modest computational resources, but exhibit several limitations. First, the generated replacements are constrained to match the token length of the original sensitive span, preventing substitutions that are shorter or longer. This constraint not only limits flexibility but also introduces a vulnerability, as adversaries may exploit this token-length consistency to infer the original content. Moreover, there is no mechanism to ensure that the replacement remains truthful to the original PII span.
@@
Larger, instruction fine-tuned LLMs such as ChatGPT [63] have also been proposed for protecting sensitive information. Instead of fine-tuning, a common way to get these models to perform a task is via an instruction (prompt) and optional examples. Such prompting methods are also integrated in existing de-identification tools, such as Microsoft Presidio [39]. Given a text, the PII detected with the Presidio pipeline can be replaced with GPT 3.5-generated alternatives in a single step [^2] See https://github.com/microsoft/presidio/blob/main/docs/samples/python/synth_data_with_openai.ipynb. Concretely, the LLMs receives the document to be processed with the sensitive spans already replaced by their NE type (e.g., PERSON or LOCATION), and is prompted to generate replacements of the same type. This makes it possible to use a commercial, cloud-based LLM without having to send the original document to third parties. Nevertheless, this method requires complete confidence in the detection step, with major privacy implications if any PII is missed. Furthermore, like for the aforementioned BERT-based approaches, truthfulness is neglected, with the generated replacements only restricted to be of the same NE type as the original span.
@@
The only LLM-based approach explicitly focused on generating truthful replacements (that is, replacements that subsume the semantic content of the original text span) is the one described in [18]. This study targets the task closely related to sanitization of abstracting text spans containing self-disclosure, “the communication of personal information to others”. The categories of personal information employed in this study are a fine-grained version of those typically considered by previous anonymization approaches, and mostly fall under demographic attributes. The sensitive text spans in [18], however, are often phrases and clauses (written in first person) rather than words or named entities. Their proposed task can be thus viewed as a specialized case of protecting sensitive information. PII protection is performed step-wise, where self-disclosure spans are first detected with a BERT model fine-tuned on a manually annotated dataset. The text spans expressing personal information are then re-written with an open-weights instruct LLM (Llama-2-7B) fine-tuned on a dataset with replacements automatically generated using GPT-3.5 and GPT-4.
@@
In contrast, the work presented in this paper only requires access to an open-weights instruct LLM, without additional fine-tuning, to iteratively select truthful generalizations from a set of alternatives with different abstraction levels. In addition, the proposed method performs inference attacks to test whether a given replacement effectively conceals the initial content of the span.
Table 1 summarizes key characteristics of representative LLM-based sanitization methods, highlighting the underlying models, their methodologies, whether they ensure truthfulness in the replacements, and any notable limitations or contributions.
Method Approach Truthful? Model Other details
MAPA toolkit [6] and Albanese (2023) [2] Masking and sampling from top token predictions BERT (with possible fine-tuning) Restricted to same-length replacements. Vulnerable to inference attacks
Presidio’s synthetic replacements [39] Prompting with sensitive spans replaced by NE types GPT-3.5-turbo Uses closed-weights LLM. No semantic preservation guarantees
Dou et al. (2024) [18] Fine-tuning and prompting for abstraction Fine-tuned LLaMA-2-7B Targets self-disclosure. Training data partially based on GPT-3.5/4. Uses open-weights LLM
INTACT (this work) Truth-preserving replacements + inference attacks Mistral-7B-Instruct-v0.2 No fine-tuning. Evaluation of privacy risks via inference attacks. Uses open-weights LLM
Table 1: Comparison of language model-based sanitization methods.
@@ 2.3 Evaluation
A common evaluation strategy for text anonymization methods is to compare their editing choices against manual annotations, which are considered as ground truths. The anonymization quality is quantified via precision and recall metrics, which are used as proxies of the preserved utility and the attained privacy protection, respectively [36, 51, 26, 32]. Modified versions of the precision and recall metrics specifically tailored for text anonymization have also been proposed [46].
@@
However, these evaluation metrics are only applicable to assess the quality of the detection step, and are unable to assess the suitability of the replacements used for sanitization. Indeed, evaluating sanitized texts against a manually constructed ground truth is challenging because i) there can be multiple valid (i.e., non-disclosive) combinations of replacements for a given text and, ii) replacement candidates are, quite often, unbounded. Establishing a single ground truth for the evaluation is thus challenging. Moreover, requesting human annotators to provide privacy-preserving replacements (or, at the very least, chose appropriate replacements for predefined lists of candidates) requires even more effort than the –already quite costly– detection of disclosive text spans [43]. As a result, the sanitization strategies currently proposed in the literature remain largely untested [14, 31, 3, 53, 54, 52, 27, 43].